# An alternating low-rank projection approach for partial differential equations with random inputs

### v1

@@ 1 Introduction
Designing efficient solvers is a crucial and challenging problem when dealing with the stochastic Galerkin methods, as it often leads to large coupled linear systems. Since the stochastic Galerkin methods typically result in linear systems formulated in the Kronecker form [13, 14, 15], various iterative algorithms that exploit the Kronecker structure of linear systems to reduce computational costs have been extensively studied in recent years [16, 17, 18, 19, 20]. Additionally, it has been shown that the desired stochastic Galerkin solution can be approximated through low-rank approximation, which further reduces computational costs [21].
@@ 2.2 Stochastic Galerkin method
$L^{2}(D)$ ::=:= $\left\{\psi(\bm{x}):D\to\mathbf{R}\ \bigg{|}\ \int_{D}\psi^{*}(\bm{x})\psi(\bm{x})\,\mathrm{d}\bm{x}<\infty\right\},$
$L_{\rho}^{2}(\Gamma)$ ::=:= $\left\{\phi(\bm{\xi}):\Gamma\to\mathbf{R}\ \bigg{|}\ \int_{\Gamma}\rho(\bm{\xi})\phi^{*}(\bm{\xi})\phi(\bm{\xi})\,\mathrm{d}\bm{\xi}<\infty\right\},$
@@
$W_{h}^{p}:=X_{h}\otimes S_{p}:={\rm span}\left\{\psi_{s}(\bm{x})\phi_{j}(\bm{\xi})\left|\psi_{s}\in X_{h},\phi_{j}\in S_{p}\right.\right\},$
@@
$\begin{split}X_{h}={\rm span}\left\{\psi_{s}(\bm{x})\right\}_{s=1}^{n_{x}}\subseteq X(D),\ S_{p}={\rm span}\left\{\phi_{j}(\bm{\xi})\right\}_{j=1}^{n_{\xi}}\subseteq L^{2}_{\rho}(\Gamma).\end{split}$
@@ 3.2 A simultaneous low-rank projection approach
It is easy to verify that the error of rank $k$ approximation of $u^{p}(\bm{x},\bm{\xi})$ is given by
$\left\|u^{p}(\bm{x},\bm{\xi})-u^{p}_{k}(\bm{x},\bm{\xi})\right\|_{L^{2}\otimes L^{2}_{\rho}}=\left(\sum_{j=k+1}^{n_{\xi}}{\sigma}_{j}^{2}\right)^{1/2}.$
@@
as the basis for the physical and stochastic subspaces respectively, we may result in a much smaller linear system compared to (6). However, since the $D\times n_{\xi}$ quasimatrix $\mathbb{U}^{p}_{D\times n_{\xi}}$ cannot be determined in advance, some adjustments are necessary to adapt this idea to design a solver for (1).
It is worth noting that the accuracy of the rank $k$ approximation of $u^{p}(\bm{x},\bm{\xi})$ depends only on the quasimatrix $\mathbb{U}^{p}_{D\times n_{\xi}}$. In practical implementations, quadrature formulas are typically used to compute the integral in the Gram–Schmidt orthogonalization process for the QR factorization of a quasimatrix. This inspires us to construct the basis functions of the stochastic space and to determine the rank $k$ as follows:
@@
we note that $\widehat{\mathbb{S}}^{*}$ is a $\Gamma\times n_{\xi}$ quasimatrix with orthonormal rows. Denoting the rows of $\widehat{\mathbb{S}}^{*}$ by $\widehat{\phi}_{1}(\bm{\xi}),\ldots,\widehat{\phi}_{n_{\xi}}(\bm{\xi})$, the low-rank basis functions of stochastic space is given by:
$S^{(k)}=\{\widehat{\phi}_{1},\ldots,\widehat{\phi}_{k}\},$
@@
For the low-rank basis functions of the physical space, choosing the columns of the quasimatrix $\widehat{\mathbb{U}}$ becomes less effective when the mesh grid is refined. Instead, we construct the low-rank basis functions of the physical space based on Theorem 3.3. This theorem asserts that if the number of snapshots (i.e., $u^{p}(\bm{x},\bm{\xi}_{i}),\ i=1,\ldots,n$) is sufficiently large, the non-zero singular values of ${\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}$ can be arbitrarily close to the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$ with a scale factor. In other words, the range of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ aligns with the range of $\mathbb{U}^{p}_{D\times n_{\xi}}$. This suggests that the left singular vectors of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ can be used in place of the left singular vectors of $\mathbb{U}^{p}_{D\times n_{\xi}}$ as the basis functions of the physical space.
@@
$X^{(k)}=\{\widetilde{u}_{1},\widetilde{u}_{2},\ldots,\widetilde{u}_{k}\}$
as the basis functions of the physical space, where $k$ is the same number as in (18), and $\widetilde{u}_{1},\widetilde{u}_{2},\ldots,\widetilde{u}_{k}$ are the columns of $\widetilde{\mathbb{U}}$.
@@
$\widetilde{\mathbb{U}}=\mathbb{X}\bm{W}^{(k)},\ \widetilde{\mathbb{S}}^{*}=(\bm{V}^{(k)})^{*}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*},$
@@
and the matrix $\bm{V}^{(k)}$ consists of the first $k$ columns of $\widehat{\bm{V}}$ defined in (16). The simultaneous low-rank projected matrix equation is then given by
@@
where
$\bm{A}_{i}^{(k)}$ $=(\bm{W}^{(k)})^{*}\bm{A}_{i}\bm{W}^{(k)},$ $\bm{f}^{(k)}$ $=(\bm{W}^{(k)})^{*}\bm{f},$ (20)
$\bm{G}_{i}^{(k)}$ $=(\bm{V}^{(k)})^{\mathrm{T}}\bm{G}_{i}\left((\bm{V}^{(k)})^{*}\right)^{\mathrm{T}},$ $\bm{h}^{(k)}$ $=(\bm{V}^{(k)})^{\mathrm{T}}\bm{h}.$ (21)
The low-rank projection approximation of the solution can then be given by
@@
5:$\ k$longleftarrow the smallest number satisfying (18), where $\widehat{\sigma}_{1},\ldots,\widehat{\sigma}_{n_{\xi}}$ are the diagonal of $\widehat{\bm{\Sigma}}$.
6:$\bm{V}^{(k)}$longleftarrow the first $k$ columns of $\widehat{\bm{V}}$.
7:$\ \bm{\theta}$longleftarrow the most important $k$ samples selected from $\bm{\Theta}$.
8:$\bm{U}^{{\mathrm{f}}}$longleftarrow snapshots on the fine physical grid corresponding to $\bm{\theta}$.
9:$\bm{W}^{(k)}$longleftarrow the first $k$ left singular vectors of $\bm{U}^{{\mathrm{f}}}$.
10:$\bm{X}^{(k)}$$\longleftarrow$ solve (19).
@@ 3.3 An alternating low-rank projection approach
The idea is quite straightforward. After applying SimLRP, we alternatively implement the low-rank projection in the stochastic and physical spaces to refine and update the low-rank approximation. Suppose we have obtained $\bm{W}^{(k)}$, $\bm{X}^{(k)}$, $\bm{V}^{(k)}$ from the SimLRP approach. We then solve the following systems of equations in sequence:
$\sum_{i=1}^{K}\bm{A}_{i}\bm{Y}^{(k)}(\bm{G}_{i}^{(k)})^{\mathrm{T}}=\bm{f}(\bm{h}^{(k)})^{\mathrm{T}},$ (23)
$\sum_{i=1}^{K}\bm{A}_{i}^{(k)}\bm{Z}^{(k)}(\bm{G}_{i})^{\mathrm{T}}=\bm{f}^{(k)}(\bm{h})^{\mathrm{T}}.$ (24)
Note that once $\bm{Y}^{(k)}$ or $\bm{Z}^{(k)}$ is solved, $\bm{W}^{(k)}$ or $\bm{V}^{(k)}$, and thus $\bm{A}^{(k)}_{i}$ or $\bm{G}^{(k)}_{i}$, should be updated immediately. To reduce the number of iterations in solving (23) and (24), the solution from the previous step can be used to design initial guesses.
@@
5:     $\bm{Y}^{(k)}$longleftarrow solve (23) with initial guess $\bm{W}^{(k)}\bm{\Sigma}^{(k)}$.
@@
8:     $\bm{Z}^{(k)}$longleftarrow solve (24) with initial guess $\bm{\Sigma}^{(k)}(\bm{V}^{(k)})^{*}$.
@@
12:$\bm{X}^{(k)}$$\longleftarrow$ solve (19) with initial guess $\bm{\Sigma}^{(k)}$.
@@
1:function [$\bm{W}^{(k)}$,$\bm{\Sigma}^{(k)}$, $\bm{V}^{(k)}$] = TruncatedSVD($\bm{X}$,$\bm{W}^{(k)}$, $\bm{V}^{(k)}$)
@@
3:     $\bm{W}^{(k)}_{{\mathrm{temp}}}$longleftarrow first $k$ columns of $\bm{W}$.
4:     $\bm{V}^{(k)}_{{\mathrm{temp}}}$longleftarrow first $k$ rows of $\bm{V}$.
5:     $\bm{\Sigma}^{(k)}_{{\mathrm{temp}}}$longleftarrow first $k\times k$ leading principal submatrix of $\bm{\Sigma}$.

### v2

@@ 1 Introduction
Designing efficient solvers is a crucial and challenging problem when dealing with the stochastic Galerkin methods, as it often leads to large coupled linear systems. Since the stochastic Galerkin methods typically result in linear systems formulated in the Kronecker form [13, 14, 15], various iterative algorithms that exploit the Kronecker structure of linear systems to reduce computational costs have been extensively studied in recent years [16, 17, 18, 19, 20, 21, 22]. Additionally, it has been shown that the desired stochastic Galerkin solution can be approximated through low-rank approximation, which further reduces computational costs [23].
@@
We note that there are various methods developed to exploit the inherent low-rank structure of solutions to stochastic PDEs. These include approaches that incorporate tensor-based representations of the input data and operators [27, 28], as well as low-rank iterative solvers for the linear systems arising from stochastic Galerkin methods [16, 17, 20, 21]. In the former case, a tensor train format is constructed for the coefficient random field involved in the stochastic PDEs, enabling efficient assembly of the stochastic Galerkin matrix in the tensor train format with the same ranks [28]. In this work, we adopt a truncated Karhunen–Loéve (KL) expansion [6, 29] to represent the random input field and focus on the standard stochastic Galerkin method, without relying on tensorized representations of the coefficient or operator.
A key distinction between the proposed approach and existing low-rank iterative solvers lies in how the rank is handled throughout the computation. Most existing solvers operate directly on the full linear systems arising from stochastic Galerkin methods and apply low-rank techniques during the iterations. These methods typically either introduce truncation operators at each iteration to control rank growth [16, 19], or adaptively increase the rank to achieve a desired level of accuracy [17, 21]. In contrast, the methods proposed in this work determine the rank by analyzing the singular values of quasimatrices associated with the truncated gPC expansion of the solution on a coarse physical grid. This pre-analysis enables the construction of a fixed-rank approximation, thereby avoiding the need for dynamic truncation or rank adaptation during iterative procedures for solving the linear systems, such as preconditioned conjugate gradient (PCG).
Another significant difference lies in the formulation of the linear system. Instead of solving the original linear system directly, we first construct a much smaller low-rank projected linear system based on the identified low-rank subspace. Once this reduced system is solved, an approximation to the original solution can be efficiently reconstructed. As a result, the proposed method significantly reduces the computational cost compared to existing approaches.
@@ 2.2 Stochastic Galerkin method
$L^{2}(D)$ $:=$ $\left\{\psi(\bm{x}):D\to\mathbf{R}\ \bigg{|}\ \int_{D}\psi^{*}(\bm{x})\psi(\bm{x})\,\mathrm{d}\bm{x}<\infty\right\},$
$L_{\rho}^{2}(\Gamma)$ $:=$ $\left\{\phi(\bm{\xi}):\Gamma\to\mathbf{R}\ \bigg{|}\ \int_{\Gamma}\rho(\bm{\xi})\phi^{*}(\bm{\xi})\phi(\bm{\xi})\,\mathrm{d}\bm{\xi}<\infty\right\},$
@@
[…]
@@
$X_{h}=\mathrm{span}\left\{\psi_{s}(\bm{x})\right\}_{s=1}^{n_{x}}\subseteq X(D),\ S_{p}=\mathrm{span}\left\{\phi_{j}(\bm{\xi})\right\}_{j=1}^{n_{\xi}}\subseteq L^{2}_{\rho}(\Gamma).$ (5)
@@ 3.2 A simultaneous low-rank projection approach
It is easy to verify that the relative error of rank $k$ approximation of $u^{p}(\bm{x},\bm{\xi})$ is given by
$\frac{\left\|u^{p}(\bm{x},\bm{\xi})-u^{p}_{k}(\bm{x},\bm{\xi})\right\|_{L^{2}\otimes L^{2}_{\rho}}}{\left\|u^{p}(\bm{x},\bm{\xi})\right\|_{L^{2}\otimes L^{2}_{\rho}}}=\frac{\left(\sum_{j=k+1}^{n_{\xi}}{\sigma}_{j}^{2}\right)^{1/2}}{\left(\sum_{j=1}^{n_{\xi}}\sigma_{j}^{2}\right)^{1/2}}.$ (17)
@@
as the low-rank approximation subsapces of the physical and stochastic spaces respectively, we may result in a much smaller linear system compared to (7). However, since the $D\times n_{\xi}$ quasimatrix $\mathbb{U}^{p}_{D\times n_{\xi}}$ cannot be determined in advance, some adjustments are necessary to adapt this idea to design a solver for (1).
It is worth noting that the accuracy of the rank $k$ approximation of $u^{p}(\bm{x},\bm{\xi})$ depends only on the quasimatrix $\mathbb{U}^{p}_{D\times n_{\xi}}$. In practical implementations, quadrature formulas are typically used to compute the integral in the Gram–Schmidt orthogonalization process for the QR factorization of a quasimatrix. This inspires us to construct the low-rank basis functions of the stochastic space (i.e., basis functions of $S^{(k)}$) and to determine the rank $k$ as follows:
@@
we note that $\widehat{\mathbb{S}}^{*}$ is a $\Gamma\times n_{\xi}$ quasimatrix with orthonormal rows. Denoting the rows of $\widehat{\mathbb{S}}^{*}$ by $\widehat{\phi}_{1}(\bm{\xi}),\ldots,\widehat{\phi}_{n_{\xi}}(\bm{\xi})$, the low-rank approximation subspace of stochastic space is given by:
$S^{(k)}=\mathrm{span}\{\widehat{\phi}_{1},\ldots,\widehat{\phi}_{k}\},$
@@
For the low-rank basis functions of the physical space (i.e., basis functions of $X^{(k)}$), choosing the columns of the quasimatrix $\widehat{\mathbb{U}}$ becomes less effective when the mesh grid is refined. Instead, we construct the low-rank basis functions of the physical space based on Theorem 3.3. This theorem asserts that if the number of snapshots (i.e., $u^{p}(\bm{x},\bm{\xi}_{i}),\ i=1,\ldots,n$) is sufficiently large, the non-zero singular values of ${\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}$ can be arbitrarily close to the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$ with a scale factor. In other words, the range of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ aligns with the range of $\mathbb{U}^{p}_{D\times n_{\xi}}$. This suggests that the left singular vectors of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ can be used in place of the left singular vectors of $\mathbb{U}^{p}_{D\times n_{\xi}}$ as the basis functions of the physical space.
@@
$X^{(k)}=\mathrm{span}\{\widetilde{u}_{1},\widetilde{u}_{2},\ldots,\widetilde{u}_{k}\}$
as the low-rank approximation subspace of the physical space, where $k$ is the same number as in (20), and $\widetilde{u}_{1},\widetilde{u}_{2},\ldots,\widetilde{u}_{k}$ are the columns of $\widetilde{\mathbb{U}}$.
@@
$\widetilde{\mathbb{U}}=\mathbb{X}\bm{W}^{(k)},\ \widetilde{\mathbb{S}}^{*}=(\bm{V}^{(k)})^{*}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*},$ (21)
@@
and the matrix $\bm{V}^{(k)}$ consists of the first $k$ columns of $\widehat{\bm{V}}$ defined in (18). Note that the columns of the quasimatrix $\mathbb{X}$ correspond to the physical basis functions. Consequently, the columns of $\bm{W}^{(k)}$ represent the coordinates of $\widetilde{u}_{1},\widetilde{u}_{2},\ldots\widetilde{u}_{k}$ with respect to the physical basis functions $\{\psi_{s}(\bm{x})\}_{s=1}^{n_{x}}$. Similarly, the rows of the quasimatrix $(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}$ correspond to the stochastic basis functions, and the rows of $(\bm{V}^{(k)})^{*}$ represent the coordinates of $\widehat{\phi}_{1},\widehat{\phi}_{2},\ldots,\widehat{\phi}_{k}$ with respect to the stochastic basis functions $\{\phi_{s}(\bm{x})\}_{s=1}^{n_{\xi}}$.
If we use
[…]
as the approximation subspace, where
$X^{(k)}=\mathrm{span}\{\widetilde{u}_{1},\widetilde{u}_{2},\ldots,\widetilde{u}_{k}\},\ S^{(k)}=\mathrm{span}\{\widehat{\phi}_{1},\ldots,\widehat{\phi}_{k}\},$ (22)
we obtain a significantly smaller linear system compared to the full system (7). We refer to this reduced system as the low-rank projected linear system, as it results from projecting the original system (7) onto the low-rank subspace $W^{(k)}$.
By substituting the low-rank projected approximation of the solution, i.e.,
$u^{p}_{k}(\bm{x},\bm{\xi}):=\sum_{j=1}^{k}\sum_{s=1}^{k}x_{sj}^{(k)}\widetilde{u}_{s}(\bm{x})\widehat{\phi}_{j}(\bm{\xi}),$ (23)
into the weak formulation (4), and using the coordinate representation (21), we obtain the low-rank projected linear system in matrix form:
@@
where $\bm{X}^{(k)}$ collects the unknown coefficients $x_{sj}^{(k)}$ from (23), with entries defined by
$\bm{X}^{(k)}(s,j)=x_{sj}^{(k)}.$
The reduced matrices and vectors in (24) are given by:
$\bm{A}_{i}^{(k)}$ $=(\bm{W}^{(k)})^{*}\bm{A}_{i}\bm{W}^{(k)},\$$\bm{f}^{(k)}$ $=(\bm{W}^{(k)})^{*}\bm{f},$ (25)
$\bm{G}_{i}^{(k)}$ $=(\bm{V}^{(k)})^{\mathrm{T}}\bm{G}_{i}\left((\bm{V}^{(k)})^{*}\right)^{\mathrm{T}},\$$\bm{h}^{(k)}$ $=(\bm{V}^{(k)})^{\mathrm{T}}\bm{h},$ (26)
where the matrices and vectors on the right-hand sides of (25)–(26) are defined in (8)–(9) and (21).
Note that the low-rank projected matrix equation (24) involves only $k\times k$ unknowns, and can therefore be solved much more efficiently than the full matrix equation (10). Once $\bm{X}^{(k)}$ is solved from (24), the low-rank approximation of the solution can be reconstructed as
@@
On the other hand, the full solution obtained from the standard stochastic Galerkin method, as given in equation (6), can be written in the form
$u^{{\scriptsize\mbox{ap}}}(\bm{x},\bm{\xi})=\mathbb{X}\bm{U}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}.$ (28)
Since both (27) and (28) approximate the solution to problem (1), the matrix product $\bm{W}^{(k)}\bm{X}^{(k)}(\bm{V}^{(k)})^{*}$ can be viewed as a low-rank approximation to the full coefficient matrix $\bm{U}$.
@@
5:k\ k $\longleftarrow$ the smallest number satisfying (20), where $\widehat{\sigma}_{1},\ldots,\widehat{\sigma}_{n_{\xi}}$ are the diagonal of $\widehat{\bm{\Sigma}}$.
[…]
@@ 3.3 An alternating low-rank projection approach
The idea is quite straightforward. After applying SimLRP, we alternatively implement the low-rank projection in the stochastic and physical spaces to refine and update the low-rank approximation. Suppose we have obtained $\bm{W}^{(k)}$, $\bm{X}^{(k)}$, $\bm{V}^{(k)}$ from the SimLRP approach. We first update $\bm{W}^{(k)}$ and $\bm{V}^{(k)}$ based on the approximate solution obtained from (27) in the SimLRP approach. We then use
[…]
as the low-rank approximation subspace, where $X_{h}$ and $S^{(k)}$ are defined in (5) and (22) respectively. Similar to the SimLRP approach, this leads to a low-rank projected linear system in matrix form:
$\sum_{i=1}^{K}\bm{A}_{i}\bm{Y}^{(k)}(\bm{G}_{i}^{(k)})^{\mathrm{T}}=\bm{f}(\bm{h}^{(k)})^{\mathrm{T}},\\$ (30)
where $\bm{Y}^{(k)}$ collects the unknown coefficients, and the other matrices and vectors are defined in (9) and (26).
Once $\bm{Y}^{(k)}$ is solved from (30), the low-rank approximation solution, similar to the SimLRP approach, is given by
$\mathbb{X}\bm{Y}^{(k)}\widetilde{\mathbb{S}}^{*}=\mathbb{X}\bm{Y}^{(k)}(\bm{V}^{(k)})^{*}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}.$ (31)
We then update $\bm{W}^{(k)}$ and $\bm{V}^{(k)}$ based on the approximation solution obtained from (31), and use
[…]
as the low-rank approximation subspace, where $X^{(k)}$ and $S_{p}$ are defined in (22) and (5), respectively. Similar to the SimLRP approach, this leads to a low-rank projected linear system in matrix form:
$\sum_{i=1}^{K}\bm{A}_{i}^{(k)}\bm{Z}^{(k)}\bm{G}_{i}^{\mathrm{T}}=\bm{f}^{(k)}\bm{h}^{\mathrm{T}}.$ (32)
where $\bm{Z}^{(k)}$ collects the unknown coefficients, and the other matrices and vectors are defined in (8) and (25).
Once $\bm{Z}^{(k)}$ is solved from (32), the low-rank approximation solution, similar to the SimLRP approach, is given by
$\widetilde{\mathbb{U}}\bm{Z}^{(k)}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}=\mathbb{X}\bm{W}^{(k)}\bm{Z}^{(k)}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}.$ (33)
We can then update $\bm{W}^{(k)}$ and $\bm{V}^{(k)}$ based on the approximate solution obtained from (33), and use $W_{\mathrm{S}}^{(k)}$ defined in (29) as the low-rank approximation subspace.
By repeating this procedure, we obtain the alternating low-rank projection (AltLRP) approach. The method is named as such because it alternates between applying low-rank projections in the stochastic and physical spaces. To reduce the number of iterations in solving (30) and (32), the solution from the previous step can be used as an initial guess.
@@
5:  $\bm{Y}^{(k)}$ $\longleftarrow$ solve (30) with initial guess $\bm{W}^{(k)}\bm{\Sigma}^{(k)}$.
@@
8:  $\bm{Z}^{(k)}$ $\longleftarrow$ solve (32) with initial guess $\bm{\Sigma}^{(k)}(\bm{V}^{(k)})^{*}$.
@@
[…]
@@
1:function $\bm{W}^{(k)}$$\bm{\Sigma}^{(k)}$, $\bm{V}^{(k)}$] = TruncatedSVD($\bm{X}$$\bm{W}^{(k)}$, $\bm{V}^{(k)}$)
@@
3:  $\bm{W}^{(k)}_{{\mathrm{temp}}} \longleftarrow$ first $k$ columns of $\bm{W}$.
4:  $\bm{V}^{(k)}_{{\mathrm{temp}}} \longleftarrow$ first $k$ rows of $\bm{V}$.
5:  $\bm{\Sigma}^{(k)}_{{\mathrm{temp}}} \longleftarrow$ first $k\times k$ leading principal submatrix of $\bm{\Sigma}$.